OpenStudy (anonymous):

determine integral: 2x + 1/ x^2 + x dx

OpenStudy (anonymous):

u=x^2+x du=2x+1 \[\int\limits_{}^{}\frac{1}{u}du\] which is ln |x|+c ln (u)+c substitute lnx(x+1) ln(x)+ln(x+1)+c